# Autoregressive Semantic Visual Reconstruction Helps VLMs Understand Better

### v1

@@ 4.3 Visual Encoder
The visual encoder provides continuous visual features as inputs to the LMM, directly influencing the effectiveness of visual information modeling. To enhance multimodal understanding, it is crucial to employ a visual encoder that is semantically aligned with textual representations [56, 40, 55], thus enabling the extraction of high-level, semantically meaningful image features. Typically, such visual encoders adopt transformer-based [10] architecture, trained using contrastive loss [41] to align closely with textual semantics and directly convert input images into one-dimensional sequences of continuous feature vectors.
@@ 4.4 Training Recipe
As shown in Figure 2, we visualize our training recipe, which extends the standard LVLM training framework by incorporating visual supervision to enable unified autoregressive modeling over both visual inputs and textual responses. Specifically, during the pre-training stage, we focus solely on optimizing the projector and the visual head. This stage aligns visual representations sequence with the LVLM’s semantic space, allowing the model to develop an initial perception of image features by learning the mapping between continuous visual features and discrete visual token indices. In the instruction tuning stage, we further fine-tune the parameters of the LLM backbone. Leveraging diverse vision-language instruction data, the model is guided to perform deeper semantic sensing of visual content, thereby enhancing its ability to understand and reason across modalities in a more comprehensive manner.
5 Experiments
In this section, we present a comprehensive set of controlled experiments to evaluate the effectiveness of our method (ASVR) within typical LVLM’s frameworks [28] across a diverse range of multimoda understanding tasks.We begin by detailing our experimental setup. Then, we analyze the impact of different visual encoders and visual tokenizers on the model’s performance. Finally, we further validate the generalization and adaptability of our method across various LLM backbones with different parameter scales and under varying amounts of training data.
5.1 Experimental Setup
Implementation Details.
We implement our experiments baseline on the LLaVA-1.5 [27] settings only with textual supervision detaily discussed in sec 3. We utilize Vicuna-1.5-7B [65] as the LLM backbone and initialize visual encoder with the pretrained weights from SigLIP-SO400M-patch14-384 [2] to support continuous visual features for LMM. For visual tokenizer, we employ both visual appearance tokenizer and visual semantic tokenizer proposed in DualToken [43] to construct visual supervision targets, which convert input images into $27\times 27\times 8$ visual semantic or appearance token sequences, with a residual depth of $D=8$. The visual head also derived from DualToken, is integrated and aligned with the chosen visual tokenizer to ensure architectural compatibility. Additional training details and architecture of visual head are provided in Appendix. The training data is LLaVA-558K [28] and LLaVA-1.5-665K [28] for the pre-training stage and the instruction tuning stage, respectively.
Evaluation Details
We conduct a comprehensive evaluation of model’s capabilities on 14 widely used vision-language understanding benchmarks. Specifically, the general multimoda benchmarks include MMBench [33] English dev split(MMB), GQA [18], SEED-Image(SEED) [24] and MME sum [13]. For OCR-based question answering, we assessed performance on TextVQA(TVQA) [42], ChartQA(CQA) [36], DocVQA(DVQA) [37] and OCRBench(OCRB) [34] . For knowledge-based question answering, we utilize MMMU validation split [62], AI2D [21]. Additionally, we evaluated hallucination robustness on POPE [26], Hallusionbench(Hbench) [15] and visual-centric tasks on MMVP [47] and RealworldQA(RQA) [58]. Evaluation prompts can be found in Appendix.
5.2 Main Results
Table 1: The impact of ASVR under different combinations of visual tokenizers and visual encoders across multimoda understanding benchmarks. "✗" indicates the use of textual supervision only, while "✓" denotes the inclusion of visual supervision by computing additional $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$. "Sem." refers to using visual semantic tokenizer to construct visual supervision targets; "App." denotes visual appearance tokenizer; "App.+Sem." represents dual supervision, where both visual semantic and visual appearance tokenizers are used independently to compute their respective $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$, which are then summed. ASVR utilize Semantic Supervision
@@ 5.2 Main Results
Dualtoken (Discrete Visual Features)
@@
SigLIP-ViT-SO400M/14@384 (Continuous Visual Features)
@@
ASVR Sem. 59.5(+3.5) 24.3(+3.2) 35.4(+4.1) 16.4(+1.8) 66.1(+2.1) 72.8(+5.6) 66.4(+2.6) 61.5(+1.0) 33.9(+1.2) 57.0(+3.5) 54.1(+2.1) 30.0(+1.3) 33.7(+9.8) 86.3(+0.4) 51.3
@@
As shown in Table 1, with the configuration of the continuous-based visual encoder (SigLIP), we observe ASVR consistent and significant performance improvements across all 14 benchmarks, increasing the average score from 46.8 to 51.3, with 5%.
Notably, the gains are evident even on knowledge-based QA such as MMMU [62] and AI2D [21], suggesting that reconstructing and and perceiving visual inputs can enhance the model’s cognitive reasoning abilities. Furthermore, substantial improvements are also observed on fine-grained tasks such as OCRBench [34], MMVP [47], and HallusionBench [15]. In particular, HallusionBench sees an increase of nearly 10 points, further validating the effectiveness of our method. Moreover, under the configuration with a discrete-based visual encoder (DualToken), semantic visual supervision also yields notable performance gains over the baseline. This further demonstrates the generalizability and robustness of our method.
@@
We adopt SigLIP-ViT-SO400M/14@384 [63] to provide continuous visual features, while employing visual semantic tokenizer from Dualtoken [43] to generate discrete visual features; both approaches aligned with textual semantics. Our experimental results indicate that, regardless of whether autoregressive semantic visual supervision is applied, the configuration of using continuous visual features consistently outperforms its discrete features counterpart arcoss all benchmarks. This performance gap may be attributed to image feature degradation introduced by vector quantization in discrete encoding, which can lead to loss of fine-grained visual information crucial for downstream multimoda understanding.
@@ Discussion
The combination of visual encoder for provide visual features and visual semantic tokenizer for constructing semantic visual supervision targets proves to the most effective model configuration. The visual encoder avoids the visual information loss typically introduced by vector quantization, thereby providing better visual inputs for the LMM. Meanwhile, semantic supervision guides the LVLM reconstruct high-level, semantically meaningful aspects of the image, which are benefit for multimoda understanding.Notably, our findings demonstrate that continuous visual inputs with discrete semantic visual supervision targets can be seamlessly integrated into the unified autoregressive next-token prediction paradigm in the same manner as language. This formulation enables the LVLM to reconstruct and perceive visual semantic information, enhancing LVLM’s capacity for comprehensive multimoda understanding.
5.3 Method Generality
Table 2: The Generality of ASVR under different training data scale and LLM backbone across multimoda understanding benchmarks. "✗" indicates the use of textual supervision only, while "✓" denotes the inclusion of semantic visual supervision by computing additional $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$. Visual encoder(SigLIP-ViT-SO400M/14@384) are both utilized for ASVR and baseline. "/" separates the data scale used for pre-training (left) and instruction tuning (right).
OCR General Knowledge Visual-centric Hallusion
With Different Data Scale
LLaVA Vicuna-1.5-7B 2M/2M 61.6 43.8 35.4 38.7 68.4 74.9 67.9 61.7 40.6 64.6 56.1 34.8 36.9 85.6 55.1
ASVR Vicuna-1.5-7B 2M/2M 60.6(-1.0) 43.1(-0.7) 36.2(+0.8) 38.9(+0.2) 68.6(+0.2) 76.2(+1.3) 68.7(+0.8) 62.0(+0.3) 41.4(+0.8) 64.8(+0.2) 55.9(-0.2) 35.9(+1.1) 42.2(+5.3) 85.7(+0.1) 55.7
With Different LLM Backbone
LLaVA Mistral-7B 558K/665K 50.8 15.7 34.6 15.2 65.9 66.9 67.9 62.4 32.0 53.0 55.0 35.3 32.7 86.6 48.1
We validate the generalization and robustness of ASVR in enhancing multimodal understanding under different data scales and diverse LLM backbone configurations, as summarized in Table 2.
The Impact of Data Scaling
To investigate the effect of training data scale, we also evaluate ASVR under larger training data. we adopt Bunny-pretrain-LAION-2M[16] for pre-training and Bunny-v1_1-data-2M[16] for instruction tuning. We compare the performance of ASVR against the baseline across different data scales to assess its robustness and effectiveness. As shown in Table 1 and Table 2, ASVR consistently yields substantial improvements over the baseline across different training data scales. Furthermore, as the amount of training data increases, overall model performance improves. However, ASVR maintains a consistent performance margin over the baseline, demonstrating its ability to more effectively leverage additional data through autoregressive semantic visual reconstruction.
The Impact of LLM Backbone Capacities
We further evaluate the generalization capability of ASVR across different LLM backbones to examine its robustness to variations in backbone capacities and architectures. Specifically, we extend our experiments to Mistral-7B[20], which differs from the LLaMA family [1, 65]. This evaluation allows us to rigorously test the flexibility and adaptability of ASVR, assessing its performance when integrated into different LLMs.As summarized in Table 2, ASVR consistently surpasses the baseline across a variety of multimodal benchmarks, maintaining strong performance advantages regardless of backbone variations. These results demonstrating both its robustness and adaptability in diverse LLM configurations. The backbone scaling experiment will provide in Appendix.
5.4 High Resolution Adaptation
ASVR is also compatible with existing high-resolution strategies and can further enhance the multimodal understanding capabilities of LMMs. To evaluate the effectiveness of ASVR under high-resolution configurations, we upscale the input resolution of both ASVR and the baseline models to 1152 × 1152, while keeping the training conditions identical. We use LLaVA-558K[28] for the pre-training stage and LLaVA-Next-779K[30] for instruction tuning following LLaVA-Next [30].
Table 3: The High Resolution Adaptation of ASVR across multimoda understanding benchmarks. "✗" indicates the use of textual supervision only, while "✓" denotes the inclusion of semantic visual supervision by computing additional $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$. Visual encoder(SigLIP-ViT-SO400M/14@384) and 1152 × 1152 input resolution are both utilized for ASVR and baseline."/" separates the data scale used for pre-training (left) and instruction tuning (right).
@@ 5.4 High Resolution Adaptation
As shown in Table 3, under high-resolution configurations, ASVR consistently outperforms the baseline by 2% in average scores across 14 multimodal benchmarks, further demonstrating its flexibility and robustness across different input resolutions.
5.5 Ablation Study

### v2

@@ 3.3 Visual Encoder
The visual encoder provides continuous visual features as inputs to the LLM backbone, directly affecting the quality of visual modeling. To enhance multimodal understanding, it is crucial to employ a visual encoder that is semantically aligned with textual representations (wu2024vilau; qu2024tokenflow; wu2024janus), allowing the extraction of high-level and semantically meaningful image features. Typically, such visual encoders adopt transformer-based (dosovitskiy2021imageworth16x16words) architecture, trained with contrastive loss (radford2021clip) to align closely with textual semantics and directly convert input images into one-dimensional sequences of continuous feature vectors.
@@ 3.4 Training Recipe
Our training recipe is illustrated in Figure 2, which extends the standard LVLM training by incorporating visual supervision to enable unified autoregressive modeling over both visual inputs and textual responses. Specifically, during the pre-training stage, we focus solely on optimizing the projector and the visual prediction head. This stage aligns visual representations sequence with the LVLM’s semantic space, allowing the model to acquire an initial visual perception by learning the mapping from continuous visual features to discrete visual token indices. In the instruction tuning stage, we further fine-tune the LLM backbone parameters. By leveraging diverse vision-language instruction data, the model is encouraged to perform deeper semantic understanding of visual content, thereby enhancing its ability to understand and reason across modalities in a more comprehensive manner.
Table 1: Impact of ASVR with different combinations of visual encoders and visual tokenizers across multimoda understanding benchmarks. “✗” indicates training with textual supervision only, while “✓” denotes the inclusion of visual supervision via an additional $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$. “Sem.” refers to using visual semantic tokenizer to construct visual supervision targets; "App." denotes a visual appearance tokenizer; "App.+Sem." indicates dual supervision, where visual semantic and visual appearance tokenizers are used independently to compute their respective $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$, which are then summed. Our proposed ASVR utilize semantic supervision.
@@
Visual Encoder: VQ-SigLIP-ViT-SO400M/14@384 (Discrete Visual Features)
@@
Visual Encoder: SigLIP-ViT-SO400M/14@384 (Continuous Visual Features)
@@
ASVR Sem. 59.5(+3.5) 24.3(+3.2) 35.4(+4.1) 16.4(+1.8) 66.1(+2.1) 72.8(+5.6) 66.4(+2.6) 61.5(+1.0) 33.9(+1.2) 57.0(+3.5) 54.1(+2.1) 30.0(+1.3) 33.7(+9.8) 86.3(+0.4) 49.8
4 Experiments
In this section, we present a comprehensive set of controlled experiments to evaluate the effectiveness of our method (ASVR) within typical LVLM’s frameworks (liu2023llava) across a diverse range of multimoda understanding tasks. We begin by detailing our experimental setup. Then, we analyze the impact of different visual encoders and visual tokenizers on the model’s performance. Finally, we further validate the generalization and adaptability of our method across various LLM backbones with different parameter scales and under varying amounts of training data.
4.1 Experimental Setup
Implementation Details.
We implement our experiments baseline on the LLaVA-1.5 (liu2023improved) with only textual supervision as discussed in sec 2. We utilize vicuna-v1.5-7B (zheng2023judgingllmasajudgemtbenchchatbot) as the LLM backbone and initialize visual encoder with the pretrained weights from SigLIP-SO400M-patch14-384 (alabdulmohsin2023so400m) to support continuous visual features for LVLMs. For visual tokenizer, we employ both visual appearance tokenizer and visual semantic tokenizer(VQ-SigLIP)proposed in DualToken (song2025dualtoken) to construct visual supervision targets, which convert input images into $27\times 27\times 8$ visual semantic or appearance token sequences, with a residual depth of $D=8$. The visual head also derived from DualToken, is integrated and aligned with the chosen visual tokenizer to ensure architectural compatibility. Training is conducted on LLaVA-558K (liu2023llava) for pretraining and LLaVA-1.5-665K (liu2023llava) for instruction tuning.
Evaluation Details
We conduct a comprehensive evaluation of model’s capabilities on 14 widely used vision-language understanding benchmarks. Specifically, the general multimodal benchmarks include MMBench (liu2024mmbenchmultimodalmodelallaround) English dev split(MMB), GQA (hudson2019gqanewdatasetrealworld), SEED-Image(SEED) (li2023seedbenchbenchmarkingmultimodalllms) and MME sum (fu2024mmecomprehensiveevaluationbenchmark). For OCR-based question answering, we assessed performance on TextVQA(TVQA) (singh2019vqamodelsread), ChartQA(CQA) (masry2022chartqabenchmarkquestionanswering), DocVQA(DVQA) (mathew2021docvqadatasetvqadocument) and OCRBench(OCRB) (Liu_2024) . For knowledge-based question answering, we utilize MMMU validation split (yue2024mmmumassivemultidisciplinemultimodal), AI2D (kembhavi2016diagramworthdozenimages). Additionally, we evaluated hallucination robustness on POPE (li2023evaluatingobjecthallucinationlarge), Hallusionbench(Hbench) (guan2024hallusionbenchadvanceddiagnosticsuite) and visual-centric tasks on MMVP (tong2024eyeswideshutexploring) and RealworldQA(RQA) (xai2024grok). Evaluation prompts can be found in Appendix A.3.
Table 2: Generalizability of ASVR to different training data scales and LLM backbones across benchmarks. The same visual encoder (SigLIP-ViT-SO400M/14@384) is used for ASVR and the baseline. “/” separates data scales for pre-training (left), mid-training (middle, if applicable) and instruction tuning (right).
OCR General Knowledge Visual-centric Hallusion
With Different Data Scales
Baseline vicuna-v1.5-7B 2M/2M 61.6 43.8 35.4 38.7 68.4 74.9 67.9 61.7 40.6 64.6 56.1 34.8 36.9 85.6 55.1
ASVR vicuna-v1.5-7B 2M/2M 60.6(-1.0) 43.1(-0.7) 36.2(+0.8) 38.9(+0.2) 68.6(+0.2) 76.2(+1.3) 68.7(+0.8) 62.0(+0.3) 41.4(+0.8) 64.8(+0.2) 55.9(-0.2) 35.9(+1.1) 42.2(+5.3) 85.7(+0.1) 55.7
Baseline vicuna-v1.5-7B 558K/4M/4M 57.2 44.1 49.6 39.2 71.7 71.7 68.7 58.2 37.9 70.7 56.7 40.0 36.1 85.5 56.2
ASVR vicuna-v1.5-7B 558K/4M/4M 60.0(+2.8) 46.5(+2.4) 51.3(+1.7) 41.7(+2.5) 72.2(+0.5) 73.2(+1.5) 69.9(+1.2) 59.8(+1.6) 39.7(+1.8) 71.8(+1.1) 57.5(+0.8) 42.0(+2.0) 37.9(+1.8) 86.9(+1.4) 57.9
With Different LLM Backbones
Baseline Mistral-7B 558K/665K 50.8 15.7 34.6 15.2 65.9 66.9 67.9 62.4 32.0 53.0 55.0 35.3 32.7 86.6 48.1
Baseline Vicuna-v1.5-13B 558K/665k 57.2 22.1 32.4 15.1 67.1 68.9 65.6 60.4 35.6 54.9 54.8 34.0 32.9 86.8 49.1
4.2 Main Results
@@
As shown in Table 1, with the configuration of the continuous-based visual encoder (SigLIP), we observe ASVR consistent and significant performance improvements across all 14 benchmarks, increasing the average score from 46.8 to 49.8, with 3%. Notably, the gains are evident even on knowledge-based QA such as MMMU (yue2024mmmumassivemultidisciplinemultimodal) and AI2D (kembhavi2016diagramworthdozenimages), suggesting that reconstructing and and perceiving visual inputs can enhance the model’s cognitive reasoning abilities. Furthermore, substantial improvements are also observed on fine-grained tasks such as OCRBench (Liu_2024), MMVP (tong2024eyeswideshutexploring), and HallusionBench (guan2024hallusionbenchadvanceddiagnosticsuite). In particular, HallusionBench sees an increase of nearly 10 points, further validating the effectiveness of our method. Moreover, under the configuration with a discrete-based visual encoder(VQ-SigLIP), semantic visual supervision also yields notable performance gains over the baseline. This further demonstrates the generalizability and robustness of our method.
@@
We adopt SigLIP-ViT-SO400M/14@384 (zhai2023siglip) to provide continuous visual features, while employing visual semantic tokenizer VQ-SigLIP (song2025dualtoken) to generate discrete visual features; both approaches aligned with textual semantics. Our experimental results indicate that, regardless of whether autoregressive semantic visual supervision is applied, the configuration of using continuous visual features consistently outperforms its discrete features counterpart arcoss all benchmarks. This performance gap may be attributed to image feature degradation introduced by vector quantization in discrete encoding, which can lead to loss of fine-grained visual information crucial for downstream multimoda understanding. More ablations on semantic tokenizers, training strategies and visual supervision are provided in Appendix A.5.
@@ Discussion
The combination of visual encoder for provide visual features and visual semantic tokenizer for constructing semantic visual supervision targets proves to the most effective model configuration. The visual encoder avoids the visual information loss typically introduced by vector quantization, thereby providing better visual inputs for the LMM. Meanwhile, semantic supervision guides the LVLM reconstruct high-level, semantically meaningful aspects of the image, which are benefit for multimoda understanding.Notably, our findings demonstrate that continuous visual inputs with discrete semantic visual supervision targets can be seamlessly integrated into the unified autoregressive next-token prediction paradigm in the same manner as language. This formulation enables the LVLM to reconstruct and perceive visual semantic information, enhancing LVLM’s capacity for comprehensive multimoda understanding. We further demonstrate that the unified autoregressive modeling paradigm consistently surpasses its denoising-based counterpart (wang2024reconstructivevisualinstructiontuning), with results provided in the Section 4.5.
Table 3: High resolution adaptation of ASVR across multimoda understanding benchmarks. We follow LLaVA-Next liu2024llava that utilize the visual encoder(SigLIP-ViT-SO400M/14@384) and high resolution input (1152 × 1152) for ASVR and baseline.
@@
4.3 Method Generalizability
We validate the generalization and robustness of ASVR under different data scales and LLM backbone configurations, as summarized in Table 2.
The Impact of Data Scaling
To investigate the effectiveness of ASVR under varying training data scales, we follow Bunny (Bunny-pretrain-LAION-2M (he2024efficientmultimodallearningdatacentric) for pre-training and Bunny-v1_1-data-2M (he2024efficientmultimodallearningdatacentric) for instruction tuning) and LLaVA-OV (li2024llavaonevisioneasyvisualtask) (558K for pretraining, 4M for midtraining and 4M for instruction tuning follwing LLaVA-OV  (li2024llavaonevisioneasyvisualtask) training recipe). As shown in Table 1 and Table 2, ASVR consistently yields substantial improvements over the baseline across different training data scales, demonstrating its ability to effectively leverage additional data through autoregressive semantic visual reconstruction.
The Impact of LLM Backbone Capacities
We further evaluate the generalizability of ASVR across different LLM backbones. Specifically, we extend our experiments to Mistral-7B(jiang2023mistral) and Vicuna-v1.5-13B, which differ from Vicuna-v1.5-7B in model family and scale (zheng2023judgingllmasajudgemtbenchchatbot). As shown in Table 2, ASVR consistently surpasses the baseline across multimodal understanding benchmarks, maintaining strong performance advantages regardless of backbone variations. These results demonstrating both its robustness and adaptability in diverse LLM configurations. The backbone scaling experiment and clear scaling law table will provide in Appendix A.4.
4.4 High-resolution Adaptation
@@
ASVR is also compatible with existing high-resolution strategies and can further enhance the multimodal understanding capabilities of LVLMs. To evaluate the effectiveness of ASVR under high-resolution configurations, we upscale the input resolution of both ASVR and the baseline models to 1152 × 1152, while keeping the training conditions identical. We use LLaVA-558K(liu2023llava) for the pre-training stage and LLaVA-Next-779K(liu2024llava) for instruction tuning following LLaVA-Next settings (liu2024llava). As shown in Table 3, under high-resolution configurations, ASVR consistently outperforms the baseline by 2% in average scores across 14 multimodal benchmarks, further demonstrating its flexibility and robustness across different input resolutions.
4.5 Comparison with ROSS
Table 4: The detailed comparision between ASVR and ROSS ablation variant. ASVR achieves the best performance under identical training conditions. ROSS models visual information through a denoising approach, whereas ASVR adopts unified autoregressive paradigm. The SigLIP-ViT-SO400M/14@384 is utilized for semantic visual supervision and VAE features is appearance visual supervision.
Method Visual Supervision LLM backbone Visual Modeling Data TVQA DVQA OCRB CQA MMB MME SEED GQA MMMU AI2D RQA MMVP Hbench POPE AVG
ROSS Apperance VAE Vicuna-v1.5–7B Denoising LLaVA–Next 56.3 39.6 35.9 41.0 65.6 71.7 65.9 61.6 34.4 65.5 55.0 33.3 28.9 85.9 52.9
ROSS Semantic Siglip Vicuna-v1.5–7B Denoising LLaVA–Next 57.5 40.2 37.4 42.5 67.0 70.5 66.2 62.1 34.9 64.6 55.7 30.1 31.2 85.8 53.3
ASVR Semantic Siglip Vicuna-v1.5–7B Autoregressive LLaVA–Next 58.6 40.6 39.7 43.4 67.9 73.0 67.5 62.9 34.2 65.8 55.4 36.8 39.2 85.9 55.1
ROSS (wang2024ross) reconstructs continuous, appearance-level visual features (VAE features) through denoising, whereas our ASVR reconstructs discrete, semantic-level visual indices (such as discretized SigLIP features) via autoregression. We conduct experiments using the LLaVA-Next dataset (liu2024llava) under identical training settings, clearly demonstrating that the ASVR-trained model consistently outperforms the ROSS ablation variants across multiple multimodal evaluation metrics. We also implement an additional variant of ROSS that reconstructs continuous semantic-level features (SigLIP features) through denoising. The result is shown in the table below shown in Table 4.
Our ASVR-trained model achieves the best performance, indicating that autoregressive semantic visual reconstruction (ASVR) is superior to both denoising semantic visual reconstruction ablation variants and even denoising appearance visual reconstruction (ROSS) ablation variants. We attribute this performance gap to a fundamental alignment principle: LLMs are inherently trained to model high-level semantic information. Therefore, when the visual supervision is semantically aligned with textual inputs—as in ASVR—it naturally leads to better integration and understanding. In contrast, reconstructing low-level visual features (as in appearance-based ROSS) lacks semantic alignment and can even hinder comprehension. Since tasks such as VQA rely heavily on semantic reasoning, reconstructing semantic visual information is more effective for enhancing multimodal understanding.
4.6 Qualitative Comparison
Figure 3: Qualitative comparison on attention maps, where we keep the same LLM and training data. ASVR urges the model to focus on specific image contents corresponding to the question with higher attention values.
We visualize attention-score maps from several cases, illustrating the attention distribution of the last token with respect to all visual tokens, as shown in Figure 3. Compared to the baseline (LLaVA), our ASVR method consistently demonstrates more precise focus on image regions relevant to the given textual query. This highlights that incorporating semantic visual supervision via the autoregressive semantic visual reconstruction objective $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$ effectively enhance its ability to accurately associate textual descriptions with corresponding visual elements. More comparison on attention maps are shown in Appendix A.2.
5 Related Work
Large Vision Language Models
The rapid progress in large language models (LLMs)(bai2023qwenlan; llama3modelcard; touvron2023llama2; bi2024deepseek; openai2023gpt4; openai2023chatgpt) has showcased their strong generalization and remarkable instruction-following capabilities. To further expand these strengths for interpreting and interacting with the world through both visual and linguistic channels. There has been growing interest in Large Vision-Language Models (LVLMs)(liu2023llava; liu2023improved; liu2024llava), typically trained using a straightforward two-stage visual instruction tuning paradigm (liu2023llava), and align visual features extracted by visual encoder with the knowledge and reasoning capabilities of LLMs through the lightweight projector. This process involves jointly training the projector and the LLM on visual instruction datasets, with optional fine-tuning of the visual encoder. However, supervision is limited to text outputs. ASVR introduces a novel autoregressive visual semantic supervision mechanism that encourages the LVLM to reconstruct semantic visual tokens, enhancing its multimodal understanding capabilities.
Visual Autoregression for LVLMs
Recent approaches (team2024chameleon; qu2024tokenflow; wang2024emu3; wu2024vilau; wu2024janus) introduce autoregressive visual supervision via visual tokenizers, such as VQGAN (esser2021vqgan) and VQ-VAE (oord2018neuraldiscreterepresentationlearning), enabling LVLMs to support both multimodal understanding and image generation by predict relevant next visual tokens, which are then decoded into images. In contrast,  ASVR focuses specifically on enhancing the multimodal understanding capability of LVLMs. Rather than generating images,  ASVR employs autoregressive visual supervision to reconstruct semantic visual tokens within the given continuous image features as input. While prior methods are generative,  ASVR adopts the reconstructive approach aimed at promoting perception of visual information.
Reconstructive Objectives for LVLMs
ROSSwang2024reconstructivevisualinstructiontuning introduces visual supervision for LVLMs by applying denoising objective to reconstruct continuous, appearance-level visual features (VAE features). In contrast,  ASVR proposes a unified approach by employing autoregressive objective—analogous to that used for text—to reconstruct semantic visual tokens. This design enables seamless integration of visual and textual information under a unified next-token prediction paradigm.
6 Conclusion
In summary, we introduced Autoregressive Semantic Visual Reconstruction (ASVR), enabling joint learning of visual and textual modalities within a unified autoregressive framework and effectively improving multimodal understanding capability of LVLMs. ASVR explicitly integrates semantic visual supervision on visual inputs to foster deep perception. Our findings indicate that autoregressively reconstructing semantic visual representations of images consistently enhances performance across diverse multimodal tasks and also outperform its denoising-based counterpart. This effectiveness is robust across different visual feature types, LLM backbone capacities, data scales, and high-resolution scenarios, underscoring ASVR’s adaptability, scalability and versatility.
Limitations
Due to computational resource constraints, the current implementation of ASVR does not incorporate video understanding capabilities or dynamic resolution functionalities. However, these features are readily transferable and can be integrated in future iterations. In subsequent work, we plan to scale the model to support these functionalities, while also expanding its capabilities to include additional modalities. Furthermore, we aim to extend ASVR’s capacity to enable multimodal generation, thereby broadening its applicability in various domains.
Appendix A Appendix
A.1 Use of LLMs in in Paper Writing
In preparing this paper, Large Language Models (LLMs) were employed to support the refinement of writing. Their role was limited to enhancing the linguistic presentation of the paper by improving readability, clarity, and stylistic consistency. Specifically, the models were used for tasks such as rephrasing sentences, checking grammar, and streamlining the flow of the text. We emphasize that the LLMs were not involved in generating research ideas, designing methodologies, or conducting experiments. All conceptual development, methodological design, and analytical work were carried out solely by the authors. The contribution of the LLMs was restricted to language-level improvements and did not extend to the scientific substance of the work. The authors retain complete responsibility for the content of this paper, including passages revised with LLM assistance. Care has been taken to ensure that the use of LLMs complies with ethical standards and does not give rise to plagiarism or any form of scientific misconduct.
A.2 Qualitative Comparison
Figure 4: More qualitative comparison on attention maps, where we keep the same LLM and training data. With extra vision-centric supervision signals, ASVR urges the model to focus on specific image contents corresponding to the question with higher attention values.
A.3 Evaluation Prompts
All prompts used for evaluation benchmarks are released and summarized in Table5 following Cambrian-1 (tong2024cambrian1fullyopenvisioncentric).
Table 5: Listing the prompts used in the evaluation of each benchmark.
Benchmark Prompt
TextVQA (singh2019vqamodelsread) Answer the question using a single word or phrase.
DocVQA (mathew2021docvqadatasetvqadocument) Answer the question using a single word or phrase.
OCRBench (Liu_2024) Give the short answer directly.
ChartQA (masry2022chartqabenchmarkquestionanswering) Answer the question using a single number or phrase.
MMBench (liu2024mmbenchmultimodalmodelallaround) Answer with the option’s letter from the given choices directly.
MME (fu2024mmecomprehensiveevaluationbenchmark) Answer the question using a single word or phrase.
SEED-Image (li2023seedbenchbenchmarkingmultimodalllms) Answer with the option’s letter from the given choices directly.
GQA (hudson2019gqanewdatasetrealworld) Answer the question using a single word or phrase.
MMMU (yue2024mmmumassivemultidisciplinemultimodal) Answer with the option’s letter from the given choices directly.
AI2D (kembhavi2016diagramworthdozenimages) Answer with the option’s letter from the given choices directly.
RealworldQA (xai2024grok) Please answer directly with only the letter of the correct option and nothing else.
MMVP (tong2024eyeswideshutexploring) Answer with the option’s letter from the given choices directly.
Hallusionbench (guan2024hallusionbenchadvanceddiagnosticsuite) Answer the question using a single word or phrase.
POPE (li2023evaluatingobjecthallucinationlarge) Answer the question using a single word or phrase.
A.4 The Scalability of ASVR
we show the clear scaling study along two axes in Table 6 and Table 7:
Data scaling We train on four datasets—LLaVA-1.5-665K (liu2023llava), LLaVA-Next-779K (liu2024llavanext), Bunny-2M (he2024efficientmultimodallearningdatacentric), and LLaVA-OV-4M (li2024llavaonevisioneasyvisualtask) to isolate the effect of data volume.
Table 6: The scaling relationship between computational cost (FLOPs) and average performance score across different scale dataasets with the same LLM backbone-vicuna-v1.5-7B.
Data FLOPs (×1e19) Avg Score
LLaVA-1.5-665K (liu2023llava) 1.53 49.8
LLaVA-Next-779K (liu2024llavanext) 2.49 55.1
Bunny-2M (he2024efficientmultimodallearningdatacentric) 4.52 55.7
LLaVA-OV-4M (li2024llavaonevisioneasyvisualtask) 7.51 57.9
Backbone scaling Using the same Vicuna family, we vary only the parameter count and scale up to a 13B model (the maximum allowed by our computational budget).
Table 7: The scaling relationship between computational cost (FLOPs) and average performance score across different scale backbone parameters with the same training dataset-LLaVA-1.5-665k.
LLM Backbone FLOPs (×1e19) Avg Score
vicuna-v1.5-7B 1.53 49.8
vicuna-v1.5-13B 2.58 52.4
A.5 Ablation Study